### v2

@@ Abstract
In this paper, we consider a sensor placement problem where sensors can move within a network over time. Sensor placement problem aims to select $K$ sensor positions from $N$ candidates where $K<N$. Most existing methods assume that sensor positions are static, i.e., they do not move, however, many mobile sensors like drones, robots, and vehicles can change their positions over time. Moreover, underlying measurement conditions could also be changed, which are difficult to cover with statically placed sensors. We tackle the problem by allowing the sensors to change their positions in their neighbors on the network. We sequentially learn the dictionary from a pool of observed signals on the network based on graph sampling theory. Using the learned dictionary, we dynamically determine the sensor positions such that the non-observed signals on the network can be best recovered from the observations. In experiments, we validate the effectiveness of the proposed method via the mean squared error (MSE) of the reconstructed signals. The proposed dynamic sensor placement outperforms the existing static ones for both synthetic and real data.
@@ 1 INTRODUCTION
In many cases, sensors can be assumed to form a network, including wireless sensor networks, the Internet, and power grids. Therefore, sensor placement on a network, i.e., graph, has been widely studied [13]. In the graph-based sensor placement problem, $N$ nodes of the graph represent possible sensor positions, the edges correspond to relationships among the sensors, and the signal value on a node is a measurement.
@@
To overcome the limitation of the above-mentioned static sensor placement, we consider a dynamic sensor placement problem where sensors can flexibly change their positions over time. Typically, we consider a dynamic placement: sensors that can move dynamically like drones and robots.
To realize the dynamic sensor placement, we sequentially learn the dictionary for signal reconstruction from a pool of observed signals on a network (graph signals) by utilizing sparse coding. We then dynamically determine the sensor positions at every time instance such that the non-observed graph signal can be best recovered from the sampled graph signal.
@@
where $\mathbf{u}_{i}$ is the $i$th column of an orthonormal matrix $\mathbf{U}$. It is obtained by the eigendecomposition of the graph Laplacian $\mathbf{L}=\mathbf{U\Lambda U}^{\top}$ with the eigenvalue matrix $,\mathbf{\Lambda}=\mathrm{diag}\,(\lambda_{0},\lambda_{1},$$\ldots,\lambda_{N-1})$. We refer to $\lambda_{i}$ as a graph frequency.
@@ 2 GRAPH SIGNAL SAMPLING
In this section, we introduce graph signal sampling. We first review sampling theory on graphs and then describe a sampling set selection on graphs in [17].
@@ 2.1 SAMPLING THEORY ON GRAPHS
Here, we assume graph signals (i.e., sensor measurements) inherit the underlying graph information like smooth signals on a network dominant at low graph frequency [18]. Suppose that a graph signal is characterized by the following model [19]:
@@
$\mathbf{x}=\sum_{i=0}^{M-1}d[i]\mathbf{u}_{i}=\mathbf{U}_{\mathcal{VM}}\mathbf{d},$ (3)
@@
$\mathbf{x}=\sum_{i=0}^{M-1}d[i]\mathbf{1}_{\mathcal{T}_{i}}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]\mathbf{d},$ (4)
where $\mathcal{T}_{i}\,(i=1,\ldots,M)$ is the nonoverlapping subset of nodes for the $i$th cluster: $[\mathbf{1}_{\mathcal{T}_{i}}]_{n}=1$ if the node $n$ is in $\mathcal{T}_{i}$ and 00 otherwise [25]. In this case, the generator matrix is $\mathbf{A}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]$ where $M$ is the number of clusters.
@@ 2.2 D-OPTIMAL SAMPLING STRATEGY
$[\mathbf{G}]_{n,m}=$ $\sum_{i=0}^{N-1}\sum_{p=0}^{P}\alpha_{p}\lambda^{p}u_{i}[n]u_{i}[m]$
$=$ $\left[\mathbf{U}\left(\sum_{p=0}^{P}\alpha_{p}\mathbf{\Lambda}^{p}\right)\mathbf{U}^{\top}\right]_{n,m},$ (7)
where $\{\alpha_{p}\}$ are arbitrary coefficients. In graph signal processing, (7) is referred to as a $P$-hop localized filter[30], whose nonzero response is limited within $P$-hop neighbors of the target node. This implies that the coverage area of all sensors is limited to $P$-hop.
@@
where $e$ is a properly designed cost function. Typically, $e$ is designed based on the optimal experimental designs [31], such as A-, E-, and D-optimality. D-optimality is one of the popular experimental designs and is widely used in graph signal sampling. The D-optimality is more appropriate than other optimality designs if we aim to recover graph signals from non-ideally sampled signals [29].
@@
where $\mathbf{Z}=\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}$ and thus $\mathbf{Z}_{\mathcal{M}}=\mathbf{I}_{\mathcal{MV}}\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}\mathbf{I}^{\top}_{\mathcal{MV}}$. Recall that $\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S}$ is invertible when the DS condition is satisfied. Therefore, the maximization of $\det(\mathbf{Z}_{\mathcal{M}})=\det(\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S})=|\det(\mathbf{S}^{\top}\mathbf{A})|^{2}$ leads to the best recovery of graph signals.
@@ 3 DYNAMIC SENSOR PLACEMENT ON GRAPHS
In this section, we propose a dynamic sensor placement problem based on sampling theory on graphs. First, we devise online dictionary learning based on sparse coding. Second, we derive the control law of sensors. Fig. 1 shows the overview of the proposed method. The flow of the proposed method is as follows:
@@
    The dictionary is sequentially learned from previous observations (Section 33.1).
@@
    Time-varying graph signals are sampled and reconstructed based on graph sampling theory.
@@
    Sensor positions are dynamically determined based on the dictionary (Section 33.2).
@@ 3.1 ONLINE DICTIONARY LEARNING
In sampling and reconstruction introduced in Section 22.1, $\mathbf{A}$ in (2) is assumed to be fixed. Note that (10) is regarded as a static sensor placement problem since the sensor positions are determined only once. In contrast, we consider the time-varying $\mathbf{A}_{t}$ in this paper where the subscript $t$ denotes the time instance (see Table 1). In this setting, the optimal sensor positions can be changed according to $\mathbf{A}_{t}$.
@@
First, we suppose the measurement model as follows:
$\mathbf{Y}_{t}=f(\mathbf{X}_{t-1})+\mathbf{M},$ (12)
where $f$ is a mapping from $\mathbf{X}_{t-1}$ to $\mathbf{X}_{t}$, i.e., $\mathbf{X}_{t}=f(\mathbf{X}_{t-1})$, and $\mathbf{M}$ is additive white Gaussian noise. That is, every observation is generated from signals in the previous time instance. Although the mapping of the time-evolution, $f$, is generally unknown, we suppose that $f$ is Lipschitzian, i.e., there exists some constant $L$ that satisfies the following inequality.
@@
This assumption implies that signals vary smoothly in time [^1] Note that (14) is necessary to stably learn the dictionary, while we do not explicitly estimate $f$ in the following part..
Next, we consider estimating the generator $\mathbf{A}_{t}$ as a dictionary from (the estimated) $\tilde{\mathbf{X}}_{t-1}$ in (11) based on the assumption in (14) where $\tilde{\cdot}$ corresponds to the estimated vector/matrix.
Since the exact value of $M$ in (3) and (4) may not be known in general, herein, we suppose $M\ll D$, and denote a dictionary and expansion coefficient by $\mathbf{A}_{t}\in\mathbb{R}^{N\times D}$ and $\mathbf{d}_{t}\in\mathbb{R}^{D}$, respectively. Therefore, we seek the optimal $\mathbf{A}_{t}$ such that the number in nonzero values of $\mathbf{d}_{t}$ is close to $M$, i.e., $\mathbf{d}_{t}$ is sparse. This setting is similar to well-studied dictionary learning problems [33, 34].
@@
Here, we assume that the graph $\mathcal{G}$, initial generator matrix $\mathbf{A}_{0}$, and expansion coefficients matrix $\mathbf{D}_{0}$ for training are given, where $\mathbf{A}_{0}$ and $\mathbf{D}_{0}$ are arbitrary matrices. Following from (14), we can suppose that $\mathbf{X}_{t}$ is sufficiently close to $\mathbf{X}_{t-1}$. As a result, we formulate the following problem:
$\mathop{\rm arg~{}min}\limits_{\mathbf{A}_{t},\mathbf{D}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\mathbf{D}_{t}\|_{F}^{2}+\mu\|\mathbf{D}_{t}\|_{1}+\eta\|\mathbf{A}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2},$ (15)
where $\tilde{\mathbf{X}}_{t-1}=\tilde{\mathbf{A}}_{t-1}\tilde{\mathbf{D}}_{t-1}$, and $\mu$ and $\eta$ are the parameters. The first term in (15) is the data fidelity term for $\mathbf{A}_{t}$. The second and third terms are for the regularization controlling the sparsity of $\mathbf{D}_{t}$ and the temporal variation of $\mathbf{A}_{t}$, respectively.
Note that we need to solve (15) with respect to two variables $\mathbf{A}_{t}$ and $\mathbf{D}_{t}$, where they jointly form a nonconvex optimization. In this paper, we divide (15) into two independent problems with respect to $\mathbf{A}_{t}$ and $\mathbf{D}_{t}$, and solve them alternately similarly to the method in [34].
First, we solve (15) with respect to $\mathbf{A}_{t}$ by fixing $\mathbf{D}_{t}$ as $\tilde{\mathbf{D}}_{t-1}$. In this case, we easily obtain the closed-form solution for the dictionary as follows:
$\mathbf{A}_{t}$ $=\mathop{\rm arg~{}min}\limits_{\mathbf{A}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\tilde{\mathbf{D}}_{t-1}\|_{F}^{2}+\eta\|\mathbf{A}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2}$
@@
Second, we update $\mathbf{D}_{t}$ in (15) by using $\mathbf{A}_{t}$ in (16), i.e.,
$\mathbf{D}_{t}=\mathop{\rm arg~{}min}\limits_{\mathbf{D}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\mathbf{D}_{t}\|_{F}^{2}+\mu\|\mathbf{D}_{t}\|_{1}.$ (17)
@@
$\mathbf{D}_{t}=\mathop{\rm arg~{}min}\limits_{\mathbf{D}_{t}}g(\mathbf{D}_{t})+h(\mathbf{D}_{t}),$ (18)
@@
$\mathbf{D}_{t,n+1}={\rm prox}_{\gamma h}(\mathbf{D}_{t,n}-\gamma\nabla g(\mathbf{D}_{t,n})),$ (19)
where $\gamma$ is the step size, $\nabla g$ is the gradient of $g$, and ${\rm prox}_{\gamma h}$ is defined as
@@
$\mathbf{D}_{t,n+1}$ $=S_{\gamma\mu}(\mathbf{D}_{t,n}-\gamma\nabla g(\mathbf{D}_{t,n}))$
$=S_{\gamma\mu}\{\mathbf{D}_{t,n}-2\gamma\mathbf{A}_{t}^{\top}(\mathbf{A}_{t}\mathbf{D}_{t,n}-\tilde{\mathbf{X}}_{t-1})\},$ (22)
We set the step size as $\gamma\leq 1/{\lambda_{\text{max}}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})}$ according to the convergence condition of ISTA [38]. We iterates (22) until $\|\mathbf{D}_{t,n+1}-\mathbf{D}_{t,n}\|_{F}^{2}<\epsilon$ where $\epsilon$ is a small constant. We summarize the optimization of (15) in Algorithm 1.
@@
       $\mathbf{A}_{t}=(\eta\tilde{\mathbf{A}}_{t-1}+\tilde{\mathbf{X}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})(\eta\mathbf{I}_{D}+\tilde{\mathbf{D}}_{t-1}\mathbf{D}_{t-1}^{\top})^{-1}$
$\mathbf{D}_{t}=\tilde{\mathbf{D}}_{t-1}$
while $\|\mathbf{D}_{t,n}-\mathbf{D}_{t,n-1}\|_{F}^{2}\geq\epsilon$ do
       $\mathbf{D}_{t,n+1}=S_{\gamma\mu}\{\mathbf{D}_{t,n}-2\gamma\mathbf{A}_{t}^{\top}(\mathbf{A}_{t}\mathbf{D}_{t,n}-\tilde{\mathbf{X}}_{t-1})\}$

### v4

@@ Abstract
In this paper, we consider a sensor placement problem where sensors can move within a network over time. Sensor placement problem aims to select $K$ sensor positions from $N$ candidates where $K<N$. Most existing methods assume that sensor positions are static, i.e., they do not move, however, many mobile sensors like drones, robots, and vehicles can change their positions over time. Moreover, underlying measurement conditions could also be changed, which are difficult to cover with statically placed sensors. We tackle the problem by allowing the sensors to change their positions in their neighbors on the network. We dynamically determine the sensor positions based on graph signal sampling theory such that the non-observed signals on the network can be best recovered from the observations. For signal recovery, the dictionary is learned from a pool of observed signals. It is also used for the sensor position selection. In experiments, we validate the effectiveness of the proposed method via the mean squared error of the reconstructed signals. The proposed dynamic sensor placement outperforms the existing static ones for both synthetic and real data.
@@ 1 INTRODUCTION
Data on a network can be measured by sensors located at its nodes like wireless sensor networks, the Internet, and power grids [2]. Therefore, sensor placement on a network, i.e., graph, has been widely studied [13]. In the graph-based sensor placement problem, nodes of the graph represent possible sensor positions, edges correspond to relationships among the sensors, and the signal value on a node is a measurement.
@@
To overcome the limitation of the above-mentioned problem, in this paper, we propose a dynamic sensor placement where sensors can flexibly change their positions over time.
To realize the dynamic sensor placement, we sequentially learn the dictionary for signal reconstruction from a pool of observed signals on a network (i.e., graph signals) by utilizing sparse coding. We then dynamically determine the sensor positions at every time instance based on graph signal sampling theory [17], such that the non-observed graph signal can be best recovered from the sampled graph signal.
@@
where $\mathbf{u}_{i}$ is the $i$th column of an orthonormal matrix $\mathbf{U}$. It is obtained by the eigendecomposition of the graph Laplacian $\mathbf{L}=\mathbf{U\Lambda U}^{\top}$ with the eigenvalue matrix $,\mathbf{\Lambda}=\mathrm{diag}\,(\lambda_{0},\lambda_{1},$$\ldots,\lambda_{N-1})$. Without loss of generality, we assume the eigenvalues and corresponding eigenvectors are ordered in ascending order of the eigenvalues, i.e., $0=\lambda_{0}\leq\lambda_{1}\leq\cdots\leq\lambda_{N-1}$ and $\mathbf{U}=\begin{bmatrix}\mathbf{u}_{0},\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}\end{bmatrix}$. Since the eigenvector associated by lower $\lambda_{i}$ captures smoother components on the graph, we refer to $\lambda_{i}$ as a graph frequency.
@@ 2 GRAPH SIGNAL SAMPLING
In this section, we introduce graph signal sampling. We first review sampling theory on graphs and then describe a sampling set selection on graphs as presented in [18].
@@ 2.1 Sampling Theory on Graphs
A key difference between the sampling of graph signals and that of standard signals stems from the underlying signal model [17]. For example, smooth signals on a graph are characterized by small variations in signal values between adjacent nodes. Here, we assume that sensor measurements are implicitly governed by the underlying graph topology.
Suppose that a graph signal is generally represented by the following model [17]:
@@
$\mathbf{x}=\sum_{i=0}^{M-1}d_{i}\mathbf{u}_{i}=\mathbf{U}_{\mathcal{VM}}\mathbf{d},$ (3)
@@
$\mathbf{x}=\sum_{i=0}^{M-1}d_{i}\mathbf{1}_{\mathcal{T}_{i}}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]\mathbf{d},$ (4)
where $\mathcal{T}_{i}\,(i=1,\ldots,M)$ is the nonoverlapping subset of nodes for the $i$th cluster: $\mathbf{1}_{\mathcal{T}_{i}}$ is the cluster indicator vector: $[\mathbf{1}_{\mathcal{T}_{i}}]_{n}=1$ if $n\in\mathcal{T}_{i}$ and $[\mathbf{1}_{\mathcal{T}_{i}}]_{n}=0$ otherwise [24]. In this case, the generator matrix is $\mathbf{A}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]$ where $M$ is the number of clusters.
@@
$\bm{c}=\mathbf{S}^{\top}(\mathbf{x}+\mathbf{m}),$ (5)
where $m[i]\sim\mathcal{N}(0,\sigma^{2})$ is white Gaussian noise. It is best recovered by the following transform [26, 27, 28]:
@@ 2.2 D-optimal Sampling Strategy
$[\mathbf{G}]_{nm}=$ $\sum_{i=0}^{N-1}\sum_{j=0}^{J}\alpha_{j}\lambda^{j}u_{i}[n]u_{i}[m]$
$=$ $\left[\mathbf{U}\left(\sum_{j=0}^{J}\alpha_{j}\mathbf{\Lambda}^{j}\right)\mathbf{U}^{\top}\right]_{nm},$ (8)
where $\{\alpha_{j}\}$ are arbitrary coefficients. In graph signal processing, (8) is referred to as a $J$-hop localized filter[29], whose nonzero response is limited within $J$-hop neighbors of the target node. This implies that the coverage area of all sensors is limited to $J$-hop.
@@
where $e$ is a properly designed cost function.
Typically, $e$ is designed based on the optimal experimental designs [30], such as A-, E-, and D-optimality. D-optimality is one of the popular experimental designs and is widely used in graph signal sampling. The D-optimality is more appropriate than other optimality designs if we aim to recover graph signals from non-ideally sampled signals [28].
@@
where $\mathbf{Z}=\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}$ and thus $\mathbf{Z}_{\mathcal{M}}=\mathbf{I}_{\mathcal{MV}}\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}\mathbf{I}^{\top}_{\mathcal{MV}}$ (see Table 1 and Definition 1). Recall that $\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S}$ is invertible when the DS condition is satisfied. Therefore, the maximization of $\det(\mathbf{Z}_{\mathcal{M}})=\det(\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S})=|\det(\mathbf{S}^{\top}\mathbf{A})|^{2}$ leads to the best recovery of graph signals.
@@ 3 DYNAMIC SENSOR PLACEMENT ON GRAPHS
In this section, we propose a dynamic sensor placement problem based on sampling theory on graphs. First, we derive online dictionary learning based on sparse coding. Second, we consider the control law of sensors. Fig. 1 shows the overview of the proposed method. The flow of the proposed method is outlined as follows:
@@
    The dictionary is sequentially learned from previous observations (Section 3-3.1).
@@
    Time-varying graph signals are sampled and reconstructed based on graph sampling theory (Section 2).
@@
    Sensor positions are dynamically determined based on the dictionary (Section 3-3.2).
@@ 3.1 Online Dictionary Learning
In sampling and reconstruction introduced in Section 2-2.1, $\mathbf{A}$ in (2) is assumed to be fixed. Note that (11) is regarded as a static sensor placement problem since the sensor positions are determined only once. In contrast, we consider the time-varying $\mathbf{A}_{t}$ in this paper where the subscript $t$ denotes the time instance (see Table 1). In this setting, the optimal sensor positions can be changed according to $\mathbf{A}_{t}$.
@@
We assume the measurement model as follows:
$\mathbf{X}_{t}=f(\mathbf{X}_{t-1}),$ (13)
where $f$ is a mapping from $\mathbf{X}_{t-1}$ to $\mathbf{X}_{t}$. That is, every observation is generated from signals in the previous time instance. Although the mapping of the time-evolution, $f$, is generally unknown, we suppose that $f$ is Lipschitzian, i.e., there exists some constant $L$ that satisfies the following inequality.
@@
In this paper, we assume a non-expansive mapping in (13), i.e., $f$ with the Lipschizian constant $L\leq 1$, to ensure that the signals are reasonably smooth over time. This assumption is essential for learning the dictionary stably, while we do not explicitly estimate $f$ in the following part.
Next, we consider estimating the generator $\mathbf{A}_{t}$ as a dictionary from (the estimated) $\tilde{\mathbf{X}}_{t-1}$ in (12) based on the assumption in (15), where $\tilde{\cdot}$ corresponds to the estimated vector/matrix.
Since the exact value of the number of atoms, $M$ in (3) and (4) may not be known in general, herein, we suppose $M\ll D$, and denote a dictionary and expansion coefficient by $\mathbf{A}_{t}\in\mathbb{R}^{N\times D}$ and $\mathbf{d}_{t}\in\mathbb{R}^{D}$, respectively. Therefore, we seek the optimal $\mathbf{A}_{t}$ such that the number in nonzero values of $\mathbf{d}_{t}$ is close to $M$, i.e., $\mathbf{d}_{t}$ is sparse. This setting is similar to well-studied dictionary learning problems [32, 33].
@@
Here, we assume that the graph $\mathcal{G}$, training data $\mathbf{X}_{0}$, initial generator matrix $\mathbf{A}_{0}$, and expansion coefficients matrix $\mathbf{D}_{0}$ are given, where $\mathbf{A}_{0}$ and $\mathbf{D}_{0}$ are arbitrary matrices. Following from (15), we can suppose that $\mathbf{X}_{t}$ is sufficiently close to $\mathbf{X}_{t-1}$. As a result, we formulate the following problem:
$\mathop{\rm arg~{}min}\limits_{\tilde{\mathbf{A}}_{t},\tilde{\mathbf{D}}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t}\|_{F}^{2}+\mu\|\tilde{\mathbf{D}}_{t}\|_{1}+\eta\|\tilde{\mathbf{A}}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2},$ (16)
where $\tilde{\mathbf{X}}_{t-1}=\tilde{\mathbf{A}}_{t-1}\tilde{\mathbf{D}}_{t-1}$, and $\mu$ and $\eta$ are the parameters. The first term in (16) is the data fidelity term for $\tilde{\mathbf{A}}_{t}$. The second and third terms are for the regularization controlling the sparsity of $\tilde{\mathbf{D}}_{t}$ and the temporal variation of $\tilde{\mathbf{A}}_{t}$, respectively.
Note that we need to solve (16) with respect to two variables $\tilde{\mathbf{A}}_{t}$ and $\tilde{\mathbf{D}}_{t}$, which jointly form a nonconvex optimization. In this paper, we divide (16) into two independent subproblems with respect to $\tilde{\mathbf{A}}_{t}$ and $\tilde{\mathbf{D}}_{t}$, and solve them alternately, similar to the method in [33].
First, we solve (16) with respect to $\tilde{\mathbf{A}}_{t}$ by fixing $\tilde{\mathbf{D}}_{t}$ as $\tilde{\mathbf{D}}_{t-1}$. In this case, we easily obtain the closed-form solution for the dictionary as follows:
$\tilde{\mathbf{A}}_{t}$ $=\mathop{\rm arg~{}min}\limits_{\tilde{\mathbf{A}}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t-1}\|_{F}^{2}+\eta\|\tilde{\mathbf{A}}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2}$
@@
Second, we update $\tilde{\mathbf{D}}_{t}$ in (16) by using $\tilde{\mathbf{A}}_{t}$ in (17), i.e.,
$\tilde{\mathbf{D}}_{t}=\mathop{\rm arg~{}min}\limits_{\tilde{\mathbf{D}}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t}\|_{F}^{2}+\mu\|\tilde{\mathbf{D}}_{t}\|_{1}.$ (18)
@@
$\tilde{\mathbf{D}}_{t}=\mathop{\rm arg~{}min}\limits_{\tilde{\mathbf{D}}_{t}}g(\tilde{\mathbf{D}}_{t})+h(\tilde{\mathbf{D}}_{t}),$ (19)
@@
$\tilde{\mathbf{D}}_{t,n+1}={\rm prox}_{\gamma h}(\tilde{\mathbf{D}}_{t,n}-\gamma\nabla g(\tilde{\mathbf{D}}_{t,n})),$ (20)
where $n+1$ is the number of iterations, $\gamma$ is the step size, $\nabla g$ is the gradient of $g$, and ${\rm prox}_{\gamma h}$ is defined as
@@
$\tilde{\mathbf{D}}_{t,n+1}$ $=S_{\gamma\mu}(\tilde{\mathbf{D}}_{t,n}-\gamma\nabla g(\tilde{\mathbf{D}}_{t,n}))$
$=S_{\gamma\mu}\{\tilde{\mathbf{D}}_{t,n}-2\gamma\tilde{\mathbf{A}}_{t}^{\top}(\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t,n}-\tilde{\mathbf{X}}_{t-1})\},$ (23)
We set the step size as $\gamma\leq 1/{\lambda_{\text{max}}(\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t})}$ according to the convergence condition of ISTA [37]. Therefore, the proposed algorithm is guaranteed to converge. We iterate (23) until $\|\tilde{\mathbf{D}}_{t,n+1}-\tilde{\mathbf{D}}_{t,n}\|_{F}^{2}<\epsilon$ where $\epsilon$ is a small constant. The optimization of (16) is summarized in Algorithm 1.
@@
       $\tilde{\mathbf{A}}_{t}=(\eta\tilde{\mathbf{A}}_{t-1}+\tilde{\mathbf{X}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})(\eta\mathbf{I}_{D}+\tilde{\mathbf{D}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})^{-1}$
$\tilde{\mathbf{D}}_{t}=\tilde{\mathbf{D}}_{t-1}$
while $\|\tilde{\mathbf{D}}_{t,n}-\tilde{\mathbf{D}}_{t,n-1}\|_{F}^{2}\geq\epsilon$ do
       $\tilde{\mathbf{D}}_{t,n+1}=S_{\gamma\mu}\{\tilde{\mathbf{D}}_{t,n}-2\gamma\tilde{\mathbf{A}}_{t}^{\top}(\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t,n}-\tilde{\mathbf{X}}_{t-1})\}$